OpenStudy (anonymous):

Calculate the percent composition of a compound that consists of 4.6g of Sodium and 7.1g of Chlorine Could somebody please explain to me how to solve this problem?

OpenStudy (anonymous):

If you want to find percent composition by mass, solve it as follows: \[\begin{align*} \%~ \text{Na} &= \frac{\text{mass of Na}}{\text{total mass}} \times 100\\ \\&= \frac{4.6~\text{g}}{(4.6~\text{g} + 7.1~\text{g})}\times 100\end{align*}\] Follow a similar method for Cl, or, because you only have two things in the compound, you could also use: \[\%~\text{Cl} = 100\% - (\% \text{Na})\]

OpenStudy (anonymous):

Thanks for your help mate! You are a life saver